OpenStudy (anonymous):

1. Solve the equation 4x^2 -28x=0. Check your solution(s) and state the final solution set. Show all your work.

OpenStudy (greatlife44):

you may have to use the quadratic formula but there's a way out of this

OpenStudy (greatlife44):

You could factor out 4x like this \[4x(x-7) = 0 \]

OpenStudy (greatlife44):

and then solve for x, by dividing both sides by 4x \[(x-7) = \frac{ 0 }{ 4x }\] this will give you \[(x-7) = 0 \] which would then give you x = 7 and x = 0 usually the highest power in gives you a hint at how many solutions here it's x^2 or x to the second power, so we know that there will be two solutions. so x = 7 and x = 0 let's plug these back into our original equation 4x(x-7) = 0 x = 0 4(0)(0-7) = 0*7 = 0 and x = 7 4(7)(7-7) 28(0) = 0 so sometimes you don't always have to use the quadratic formula to find out your solutions that's if you can factor and we-write your quadratic expression.

OpenStudy (greatlife44):

@phebzluvsyu \[4x ^{2}-28x \] notice how these two expressions are the same. \[4x(x-7)\]

OpenStudy (greatlife44):

I could show you the other way, the LONG way of doing this and it involves the quadratic formula

OpenStudy (greatlife44):

Quadratic functions are written in this form: \[ax ^{2}+bx+c \] and this is the quadratic formula. \[-b \pm \sqrt{ b ^{2}-4ac}/2a \]

OpenStudy (greatlife44):

we have \[4x ^{2}-28x = 0 \] we compare this to \[ax ^{2}+bx+c\] so a = 4 b = 0 and c = -28

OpenStudy (greatlife44):

we then plug this into the formula here are our solutions \[\frac{ \sqrt{ (28)^{2}-4(4)(0)} }{ 2(4) }\] \[\frac{ 28-28 }{ 8 } = 0\] \[\frac{ 28*2=56 }{ 8 } = 7 \] @phebzluvsyu